OpenStudy (anonymous):

If f(x) = x minus 3, all over 3, then what is f-1(x) ? one-thirdx - 1 3x - 1 3x - 3 3x + 3 which one?

OpenStudy (anonymous):

To find inverses, first replace f(x) with y for ease of notation: y = (x-3)/3 Second change every x to a y and every y to an x: x = (y-3)/3 Third, solve for y: x = (y-3)/3 3x = y - 3 3x + 3 = y So, the inverse is: 3x + 3